The balanced combustion reaction for C6H6 is 2C6H6(l)+15O2(g)⟶12CO2(g)+6H2O(l)+6542 kJ If 6.200 g C6H6 is burned and the heat produced from the burning is added to 5691 g of water at 21 ∘ C, what is the final temperature of the water?
Answer #1

Molar mass of C6H6,

MM = 6*MM(C) + 6*MM(H)

= 6*12.01 + 6*1.008

= 78.108 g/mol

mass(C6H6)= 6.200 g

use:

number of mol of C6H6,

n = mass of C6H6/molar mass of C6H6

=(6.2 g)/(78.11 g/mol)

= 7.938*10^-2 mol

Since Δ H is negative, heat is released

when 2 mol of C6H6 reacts, heat released = 6542.0 KJ

So,

for 7.938*10^-2 mol of C6H6, heat released = 7.938*10^-2*6542.0/2 KJ

= 2.596*10^2 KJ

= 259600 J

This heat is absorbed by water.

Given:

Q = 259600 J

m = 5691 g

C = 4.184 J/g.oC

Ti = 21 oC

use:

Q = m*C*(Tf-Ti)

259600 = 5691*4.184*(Tf-21.0)

Tf -21.0 = 10.90 oC

Tf = 31.90 oC

Answer: 31.9 oC